Operation Summary for March 12, 2010 - March 15, 2010 |
The Integrated Luminosity for the period from 3/8/10 to 3/15/10 was 60.41 pb-1. Friday March 12 The day shift began with Operations monitoring the Tevatron (TeV) store 7668 (198.06E30) and the Recycler stash (112.3E10), with the Antiproton Source stacking (12.4E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Linac personnel lowered the output to 34mA. Booster personnel continued tuning through the shift. RF personnel repaired a Booster RF station (BRF15). TeV antiproton tune studies caused a slight loss in luminosity. MI personnel conducted coalescing tuning most of the afternoon. During the evening shift, Operations disabled beam to MTest at 6 PM. Operations terminated store 7668 10:41 PM and put the TeV into shot setup at 10:53 PM. There was some downtime due to problems with BRF15 and LRF5. Saturday March 13 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (6.4E10), with Pbar stacking (49.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7669 at 12:12 AM with an Average Initial Luminosity (AIL) of 336.6E30. At 2 AM, a MI quadrupole magnet (I:Q847) tripped off due to a Klixon. An EE Support person came in and reported finding a bad power supply controller. At about 5 AM, a Recycler kicker (R:KPS3A) tripped off on fluorinert level. Operations bypassed the alarm until after the next shot. LRF5 started tripping at the end of shift. During the day shift, operators investigated the LRF5 power amplifier (PA) crowbar/permanent inhibits. LRF5 would not climb above 0.7 on the gradient without sparking. Linac personnel came in and adjusted the PA input tuning. Linac beam returned at 10:41 AM. Operations resumed sending beam to Pbar, MiniBooNE, and MTest. NuMI personnel accessed MI-65 to grease the fan bearings of the target chase air system. LRF5 tripped off at 11:16 AM and station tuning resumed. The NuMI Lambertson magnet (LAM61) would not turn on due to a filter rack breaker trip. EE Support personnel came in and reset the internal breaker. The Booster Gradient Magnet Power Supply (GMPS) and sextupole power supply (B:SEXSAF) tripped off at 11:37 AM. Booster anode limit trips occurred at the same time. Operators traced the cause to the power supply of Booster CAMAC crate 10. Booster personnel came in and investigated. Linac beam returned at 12:42 PM. Booster personnel accessed their tunnel to replace RF air filters on BRF15 and 18. Booster crate 10 still had problems so Booster personnel and operators upgraded the supply to a newer style. Booster beam returned by 4:06 PM. Booster RF tripped off due to a watchdog trip at 4:49 PM. Operators traced the Anode program limit high trips back to the Booster crate 10 power supply. Operators replaced the power supply again since the voltages were drooping. The low-level RF (LLRF) curves for frequency and bias were non-existent so Booster personnel conferred with Controls personnel about the LLRF VXI. The curves came back when crate 10 voltages were restored. BRF8 tripped off at the end of shift due to PA airflow. During the evening shift, Operations reported lower beam intensities to Pbar and NuMI because of a MI power supply (I:Q847) that wasn’t running at the correct current. Operators, with the help of EE Support personnel, were able to locally set Q847 to a value that allowed beam to run. Sunday March 14 The midnight shift began with Operations monitoring store 7669 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7669 and established store 7670 at 4:53 AM with an AIL of 289.1E30. Operators prepared MI for an access to refill the kicker (KPS3A) fluorinert. Store 7670 aborted at 7:21 AM. A breaker tripped at the Kautz Road Substation due to a problem with a disconnect (S:QP3). The remainder of shift was spent recovering from this power glitch. Many cryo system devices tripped off, but there was no quench. During the day shift, High Voltage (HV) personnel investigated the Kautz Road Substation (KRS) CB95. They reported that the cause of the breaker trip was due to a flashover inside the QP3 ‘coffin’ adjacent to the switchgear at TeV sector F4. HV and EE Support personnel reconfigured the F-Sector pulsed power from KRS to allow stacking. Beam operations returned at 11:44 AM. Operations turned the TeV on at 11:47 AM. After turn-around, TeV studies followed. Booster beam positions were off. The Proton Source Specialist came in and assisted in tuning. He adjusted the extraction kickers, Noker and Notcher on the cogged events. During the evening shift, The TeV CAMAC crate power supply at TeV sector A2 started cycling off/on and caused a loss of communication. Operators manually set the low beta magnets to zero and then bypassed them. The TeV ramp tripped off when the crate power supply was turned off. After the restore, TeV went through a turn-around. When the TeV stopped at 980 GeV after the third ramp, the sequencer was at 2 instead of 1. TeV personnel set the value back and reissued the associated clock events. The TeV came down to 150 GeV and a power supply (C:D0Q6) tripped off due to a ground fault indication. EE Support personnel came in and checked out the supply; it was okay. Operators put the TeV through a dry squeeze at 7:55 PM and a wet squeeze at 8:50 PM. Operations put the TeV into shot setup and established store 7678 at 10:27 PM with an AIL of 249.7E30. Monday March 15 The midnight shift began with Operations monitoring store 7678 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam. During the midnight shift, the MI South Anode tripped on an interlock fault; it reset after several tries. A LCW leak was redeveloping on the Klystron Debuncher water skid. Shift ended as power supply experts started work on S:QP3 ‘coffin’ change out. Water group personnel started repairs on the Klystron Debuncher skid. The Plans for Monday The plans for today are stack, stash, and store. A new MTest experiment (T932) will start today after the S:QP3 change out. The Run Coordinator will allow a four-hour MI access for RF repairs on Tuesday morning, 3/16/10. NuMI will also go down Tuesday morning for work while the MINOS detector in Sudan goes down for work. |
|
More Information Comments and Suggestions |
| last modified 03/15/2010 email Fermilab |
| Security, Privacy, Legal | |